# MrSteve: Instruction-Following Agents in Minecraft with What-Where-When Memory

### v4

@@ Abstract
Significant advances have been made in developing general-purpose embodied AI in environments like Minecraft through the adoption of LLM-augmented hierarchical approaches. While these approaches, which combine high-level planners with low-level controllers, show promise, low-level controllers frequently become performance bottlenecks due to repeated failures. In this paper, we argue that the primary cause of failure in many low-level controllers is the absence of an episodic memory system. To address this, we introduce MrSteve (Memory Recall Steve-1), a novel low-level controller equipped with Place Event Memory (PEM), a form of episodic memory that captures what, where, and when information from episodes. This directly addresses the main limitation of the popular low-level controller, Steve-1. Unlike previous models that rely on short-term memory, PEM organizes spatial and event-based data, enabling efficient recall and navigation in long-horizon tasks. Additionally, we propose an Exploration Strategy and a Memory-Augmented Task Solving Framework, allowing agents to alternate between exploration and task-solving based on recalled events. Our approach significantly improves task-solving and exploration efficiency compared to existing methods. We will release our code and demos on the project page: https://sites.google.com/view/mr-steve.
@@ 1 Introduction
In this paper, we introduce an enhanced low-level controller agent, MrSteve (Memory Recall Steve-1), designed to address the limitations of Steve-1. The key innovation of MrSteve is the integration of Place Event Memory (PEM) which is the instantiation of What-Where-When Episodic Memory. While previous approaches have explored episodic memory, they primarily target high-level planners, such as building libraries of high-level skills and plans, which do not directly improve the low-level controller’s performance. We argue it is essential for the low-level controller to possess memory capabilities. To address this, PEM manages memory more effectively, surpassing the limitations of the non-scalable FIFO memory found in transformers. PEM stores spatial and event-based information, allowing the agent to hierarchically organize and retrieve details about locations and events it has previously encountered. For PEM to be fully effective, the agent must also move directly to the desired location along with the ability to modulate between exploration and goal-directed navigation-and-execution, a capability lacking in Steve-1. Therefore, we introduce the second component of MrSteve: the Exploration Strategy and Memory-Augmented Task Solving Framework. Built upon the PEM structure, this framework enables the agent to alternate between exploration—when no relevant information is stored—and task-solving by recalling past events when applicable. This is made possible and effective with our new navigation policy, VPT-Nav.
@@ 3 Method
In this section, we describe our agent, MrSteve (Memory Recall Steve-1). We begin with the problem setting, followed by step-by-step construction of our agent’s main modules.
@@
Instruction Following Policy In sparse sequential task, a naive approach is to employ Steve-1 (Lifshitz et al., 2024), an instruction-following policy $\pi_{\text{Inst}}(a_{t}|h_{t},\tau_{n})$ that generates low-level controls (mouse and keyboard) in Minecraft. Here, $h_{t}$ is a past pixel observation sequence $x_{t-128:t}$. While past observations are processed by Transformer-XL layers in Steve-1, the model is ineffective at recalling observations from a few thousand steps ago (Lampinen et al., 2021). Additionally, the Transformer’s quadratic complexity makes it significantly inefficient to process thousands of observations. This makes Steve-1 poorly suited for sparse sequential task, as it cannot recall visited places or task-relevant resources seen in the past. To address this, we propose MrSteve which stores novel events from visited places for efficient sparse sequential task-solving.
@@ 3.1 Memory Module: Construction of Place Event Memory
Specifically, each place cluster $M_{k}$ is subdivided into event clusters, denoted as $\smash{\{E_{i}^{k}\}_{i=1}^{d_{k}}}$, where each $\smash{E_{i}^{k}}$ represents the $i$-th event cluster in $k$-th place cluster, characterized by a center embedding $e_{E_{i}}^{k}$. These event clusters are newly created and updated as the place cluster accumulates a certain number of additional experience frames. For generating event clusters, DP Means algorithm (Dinari & Freifeld, 2022) is applied on video embeddings of these frames, generated by MineCLIP, and the resulting cluster centers become a center embedding of each cluster. If the cosine similarity of the cluster centers between a newly created cluster and an existing cluster is higher than threshold $c$ (we define that two event clusters are indistinct), the two clusters are merged to prevent redundancy and ensure distinct event clusters within each place cluster. When memory capacity is exceeded, the oldest frame in the largest event cluster is removed, thus the memory can retain diverse places and distinct events within each place. More details on Place Event Memory can be found in Appendix E.
@@ 4.3 Memory-Constrained Task Solving with Memory
We demonstrated in Section 4.2 that memory is essential in solving sparse sequential tasks when there is no limitation in memory capacity. However, in real-world scenarios where memory capacity is limited, memorizing visited places and novel events becomes important. In this section, we show that Place Event Memory can benefit in this scenario. To verify this, we introduce three Memory Tasks, which are Find Water, Find Zombies’ Death Spot, and Find First-Visited House. In all tasks, the agent begins with an exploration phase, followed by a task phase. In the exploration phase, the agent follows a fixed exploratory behavior. In the subsequent task phase, the agent is given a MineCLIP embedding $e_{t}=\text{Enc}_{v}(\{o_{t}\}_{n=1}^{16})$ as a task embedding instead of text instruction, where $o_{t}$ is the pixel observation seen in the exploration phase. Then, this becomes an image goal navigation task where an agent should navigate to the location of the given image.
@@ 6 Conclusion
In this paper, we introduced MrSteve (Memory Recall Steve-1), a novel low-level controller designed to address the limitations of current LLM-augmented hierarchical approaches in general-purpose embodied AI environments like Minecraft. We argued that the primary cause of failures in many low-level controllers is the absence of an episodic memory system. To overcome this, we equipped MrSteve with Place Event Memory (PEM), a form of episodic memory that captures and organizes what, where, and when information from episodes. This allows for efficient recall and navigation in long-horizon tasks, directly addressing the limitations of existing low-level controllers like Steve-1, which rely heavily on short-term memory. Additionally, we proposed an Exploration Strategy and a Memory-Augmented Task Solving Framework, enabling agents to effectively switch between exploration and task-solving based on recalled events. Our results demonstrate significant improvements in both task-solving and exploration efficiency compared to existing methods. We believe that MrSteve opens new avenues for improving low-level controllers in hierarchical planning and are releasing our code to facilitate further research in this field.
@@ Acknowledgement
This work was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) and Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation (NRF) funded by the Ministry of Science and ICT (MSIT). The authors would like to thank the members of Machine Learning and Mind Lab (MLML) for helpful comments.
@@ Appendix A Limitations of Memory System in LLM-Augmented Agents
Memory systems in agents primarily aim to retrieve robust and accurate high-level plans for long-horizon tasks (Wang et al., 2023a; b; Zhu et al., 2023; Qin et al., 2024; Li et al., 2024). Existing works use an abstracted memory that stores the succeeded task with its plans, and often with history of observations for reliable retrieval, which is helpful when the similar plan in other tasks already exist in memory. However, these types of memory systems are not well-suited for low-level controllers for the following reasons:
  • Issues with Managing Memory Recent memory systems in Minecraft, when saving successful plans or skills, the history of observations for solving the task is all stored in FIFO manner (Wang et al., 2023b) or only task-relevant frames are stored in the plan (Li et al., 2024). However, computation complexity for retrieving the experience frames that are relevant to a new task is computationally expensive or even impossible (because the latter may not store the experience frames despite the agent observed it).
  • Lack of Mechanism for Retrieving Experience frames Current memory systems store the text instructions of succeeded tasks as keys and their plans to complete those tasks as values. The retrieval process begins by matching a task query to the task keys in memory and then filtering further using the current scene’s similarity to the stored frames before retrieving the final plan. These memory systems are targetted to retrieve the succeeded plan, but they lack a mechanism for utilizing the experience frames in the memory, which could be crucial for future tasks.
@@ C.1 Environments Setting
All tasks are implemented using MineDojo (Fan et al., 2022b). We utilize MineDojo’s success checker, where the success of each task is determined based on changes in the agent’s inventory. Hence, the agent succeeds the task if the corrensponding target item is appeared in the inventory. If the agent exceeds the time limit of each task or dies before completing assigned tasks, indicating the agent fails on the task.
@@ C.2.2 Exploration & Navigation Task Details
For this task, we used the Map 1 in Figure 8 with $100\times 100$ size, surrounded by high walls. The map includes complex terrains such as mountains and a water pond, which requires robust low-level navigation policy for successful exploration. For each episode, the agent spawns in the center of the map and explores for 6,000 steps.
@@ C.2.4 Memory Tasks
Find Zombies’ Death Spot Task At the beginning, the agent sees burning zombies for the first 1K steps. Approximately 500 steps after the start of the episode, the zombies disappear, resulting in the agent observes two distinct scenes in the same location. A goal image for navigation is selected from the observation where the zombies are burning. After 1K steps, the agent starts to travel for the rest of the exploration phase. Once the exploration phase finishes, the agent returns to the place where the zombies were burning.
@@ Appendix E Place Event Memory, Place Memory, Event Memory Details
We provide the Algorithms on Write & Read operations of Place Event Memory, and other memory variants, which are Place Memory, and Event Memory. We also provide the specifications of each memory in the following section E.4.
@@ E.1 Place Event Memory Write & Read Operations
Event Cluster Details We use the DP-Means algorithm for clustering the events. DP-Means algorithm is a Bayesian non-parametric extension of the K-means algorithm based on small variance asymptotic approximation of the Dirichlet Process Mixture Model. It doesn’t require prior knowledge on the number of clusters $K$. To run this algorithm, we first set the initial number of clusters $K^{\prime}$ and cluster the data with K++ Means initialization ($K^{\prime}$ can be $1$), then DP-Means algorithm automatically re-adjust the number of clusters based on the data points and cluster penalty parameter $\delta$. Thus, DP-Means algorithm behaves similarly to K-means with the exception that a new cluster is formed whenever a data point is farther than $\delta$ away from every existing cluster centroid.
@@
Place Cluster Details We provide a detailed explanation of how place clusters are formed. Each place cluster stores the agent’s experience frames based on its 2D ground position $(x,y)$ and head direction angle yaw. The size of a place cluster and its yaw range are defined by parameters $C$ (in Minecraft block) and $W\in(0^{\circ},180^{\circ})$, respectively, which distinguish different clusters. A place cluster is centered at a specific position $(x,y)$ with a center yaw $w$. An experience frame is assigned to a place cluster if the agent’s position falls within the range $(x-C/2,x+C/2)$ for $x$ and $(y-C/2,y+C/2)$ for $y$, and the yaw angle satisfies $(w-W/2,w+W/2)$. We use relative position and yaw for center of the place cluster implying that the center of the initial place cluster has position $(0,0)$, and yaw 00.
@@ E.3 Event Memory Write & Read Operations
Discussions on MineCLIP threshold As in Event Memory’s Write Operation in Algorithm 7, clusters generated by DP Means algorithm are either merged with an exisiting event cluster, or added as new event cluster. This is determined by MineCLIP threshold $c$, which is the criterion for separating event clusters. We note that using proper value for threshold is important for Event Memory to work reasonably across the tasks. If $c$ is too small, Event Memory will cluster experience frames with only few clusters, which may not be visually distinctive. If $c$ is too large, in extreme case, Event Memory will make event cluster for each experience frame. When memory capacity is exceeded, it will randomly remove the oldest frame (since all event clusters have same size), which behaves as a FIFO Memory.
Drawback in Event Memory Event Memory removes the frame in the largest event cluster, when memory capacity is exceeded. This indicates that the memory can retain visually distinct events. However, since it does not consider position information in clustering, similar visual frames in different places can be clustered into same event cluster. This can be fatal since task-relevant frames in different places can be removed from the memory which can be crucial in upcoming tasks. This drawback is shown in Find First-Visited House task in Section 4.3, and Long-Instruction task in Section 4.4, where MrSteve-EM showed suboptimal performances.
@@ Appendix G Exploration Experiments Details
Since Steve-1 (Lifshitz et al., 2024) has no memory module, it tends to visit same place multiple times. We also observed that Steve-1 exhibits a behavior of moving straight ahead when task instruction "Go Explore" is given, often colliding with walls and then following along the wall instead of avoiding it, which harms the Map Coverage and Revisit Count.
@@ Appendix I MrSteve in Randomly Generated Maps
In this experiment, task A could involve collecting either water, beef, or wool, while task B involves gathering logs. Additionally, we introduce a sequential task named SEQ(4), which requires the agent to solve four consecutive tasks: log, water, wool, and beef. For the ABA-Random tasks, the agent was given 12K steps, and for the SEQ(4) task, 16K steps were allowed. We evaluated following agents: MrSteve, MrSteve-EM, MrSteve-PM, MrSteve-FM, PMC-MrSteve-FM, and Steve-1.
As shown in Figure 13, MrSteve and its memory variants consistently showed higher success rate than Steve-1 across all tasks. This is because when Steve-1 is finished with task B, it tries to solve the final task A from scratch making it difficult to complete all tasks in limited time. However, MrSteve could retain the experience frames of task A in memory making it efficient to solve all tasks in time. This result suggests that augmenting memory plays a crucial role in solving sequential tasks, even in randomly generated map.
@@
In this section, we study the robustness of our proposed agent MrSteve by exploring the effects of top-$k$ selection in Mode Selector, and the size of a place cluster for Place Event Memory. For this, we used one of the tasks from ABA-Sparse task in Section 4.2, which is Wool-Log-Wool task. The success rate of MrSteve with different top-$k$’s and place cluster sizes are illustrated in Figure 14. For top-$k$ experiment, we tested five $k$ values, which are $1$, $2$, $4$, $10$, and $30$. From the results, we found that using small $k$ did not largely affect the performance. For cluster size experiment, we tested five cluster sizes, which are $1$, $2$, $4$, $6$, and $12$ in Minecraft blocks. From the results, we see that using bigger cluster size slightly lowered the performance since the center embedding of the place cluster may not be good summarization vector of the place when cluster size is large, making the agent difficult to recall task-relevant frames in the memory.
@@
Table 4: Querying Time and FLOPs for each memory types and $k$ values. We report the averages and standard errors over 1K query operations.
@@
In the previous section, we discussed the naive way of combining the idea of goal-conditioned behaivor cloning and RL finetuning for training goal-conditioned navigator. In this section, we conduct an model architecture search on VPT model to find the optimal goal-conditioned navigation model. To do this, we focused on three key design choices: 1) the input location of goal embedding, 2) training parameters, and 3) different Tr-XL adaptors . For the location of goal embedding, we consider the three tactics: (1) add the goal embedding to the image embeddings and input to Tr-XL (TrXL-Cond.), (2) add the goal embedding to Tr-XL output (Head-Cond.), and (3) add the goal embedding both at the input and output of Tr-XL (Dual-Cond.). For training parameters, we use full-finetuning or finetuning only part of the model. For the Tr-XL adaptors, we consider the adaptor from Houlsby et al. (2019), and LoRA adaptor (Hu et al., 2021b). In Figure 15, we illustrate the different model architectures for goal-conditioned VPT finetuning models.
For the model search, we summarized the navigation performance (SPL and Success Rate (SR)) of different model architectures in Table 6. Interestingly, we found that adding goal embedding to image embeddings (TrXL-Cond.) showed suboptimal behavior, while adding goal embedding to output of Tr-XL (Head-Cond., Dual-Cond.) showed good performance indicating that propagating learning signal to goal embedding through Tr-XL is difficult from RL objective. We note that using LoRA adaptor showed higher performance than adaptors from Houlsby et al. (2019). In conclusion, using Head-Cond. with LoRA finetuning performed the best. Additionally, we tested one additional model that does not update Tr-XL, which showed comparable result to the best performing model. This architecture benefits for using smaller number of learning parameters and lower gradient computations. Thus, we use the VPT-Nav trained with one of two architectures, which are Head-Cond. with LoRA finetuning, and Head-Cond. with No Tr-XL Update.
@@
Table 8: Long-Horizon Planning Tasks details. The required subgoals denotes the length of shortest plan for each task. The items that the low-level controller should obtain are listed in Items to Obtain column.
@@
LLM-generated initial plans can be inaccurate, so the low-level controllers often fails to execute subgoals. For instance, we found that the initial plans frequently omit the subgoal for creating crafting table, which is prerequisite item for the tasks in Table 8 except the carpet task. To address this problem, DEPS framework introduced the replanning procedure with the descriptor and explainer modules. The descriptor module makes a text prompt representing the inventory of the agent. The explainer modules ask the LLM the reason of failure based on the text prompt from the descriptor module. Based on the explanation, the high-level planner asks the LLM to revise the plan and the low-level controller executes subgoals from the revised plan.
Additionally, bacause the LLM does not observe the environment, the order of subgoals in a plan can be suboptimal. When some subgoals share the same prerequisite so they can be executed in any order, it is more efficient to execute subgoal, which can complete faster than the other subgoals, first. This can prevent wasting time by giving up subgoals that could be completed quickly, pursuing a subgoal whose completion time is uncertain, and then going back to search for the previously quicker task again after executing the subgoal. In DEPS, this procedure called elector module is implemented with the horizon prediction module in GSB (Cai et al., 2023a). The horizon prediction module $\mu(s_{t},\tau_{t,i}^{g})$ takes the current observation $s_{t}$ and the textual subgoals $\{\tau_{t,i}^{g}\}_{i=1}^{N}$ and predicts the time to complete each subgoal. Based on the time prediction, a subgoal to be executed is sampled from the distribution as follows:

### v5

@@ Abstract
Significant advances have been made in developing general-purpose embodied AI in environments like Minecraft through the adoption of LLM-augmented hierarchical approaches. While these approaches, which combine high-level planners with low-level controllers, show promise, low-level controllers frequently become performance bottlenecks due to repeated failures. In this paper, we argue that the primary cause of failure in many low-level controllers is the absence of an episodic memory system. To address this, we introduce MrSteve (Memory Recall Steve), a novel low-level controller equipped with Place Event Memory (PEM), a form of episodic memory that captures what, where, and when information from episodes. This directly addresses the main limitation of the popular low-level controller, Steve-1. Unlike previous models that rely on short-term memory, PEM organizes spatial and event-based data, enabling efficient recall and navigation in long-horizon tasks. Additionally, we propose an Exploration Strategy and a Memory-Augmented Task Solving Framework, allowing agents to alternate between exploration and task-solving based on recalled events. Our approach significantly improves task-solving and exploration efficiency compared to existing methods. We will release our code and demos on the project page: https://sites.google.com/view/mr-steve.
@@ 1 Introduction
In this paper, we introduce an enhanced low-level controller agent, MrSteve (Memory Recall Steve), designed to address the limitations of Steve-1. The key innovation of MrSteve is the integration of Place Event Memory (PEM) which is the instantiation of What-Where-When Episodic Memory. While previous approaches have explored episodic memory, they primarily target high-level planners, such as building libraries of high-level skills and plans, which do not directly improve the low-level controller’s performance. We argue it is essential for the low-level controller to possess memory capabilities. To address this, PEM manages memory more effectively, surpassing the limitations of the non-scalable FIFO memory found in transformers. PEM stores spatial and event-based information, allowing the agent to hierarchically organize and retrieve details about locations and events it has previously encountered. For PEM to be fully effective, the agent must also move directly to the desired location along with the ability to modulate between exploration and goal-directed navigation-and-execution, a capability lacking in Steve-1. Therefore, we introduce the second component of MrSteve: the Exploration Strategy and Memory-Augmented Task Solving Framework. Built upon the PEM structure, this framework enables the agent to alternate between exploration—when no relevant information is stored—and task-solving by recalling past events when applicable. This is made possible and effective with our new navigation policy, VPT-Nav.
@@ 3 Method
In this section, we describe our agent, MrSteve (Memory Recall Steve). We begin with the problem setting, followed by step-by-step construction of our agent’s main modules.
@@
Instruction Following Policy In sparse sequential task, a naive approach is to employ Steve-1 (Lifshitz et al., 2024), an instruction-following policy $\pi_{\text{Inst}}(a_{t}|h_{t},\tau_{n})$ that generates low-level controls (mouse and keyboard) in Minecraft. Here, $h_{t}$ is a past pixel observation sequence $i_{t-128:t}$. While past observations are processed by Transformer-XL layers in Steve-1, the model is ineffective at recalling observations from a few thousand steps ago (Lampinen et al., 2021). Additionally, the Transformer’s quadratic complexity makes it significantly inefficient to process thousands of observations. This makes Steve-1 poorly suited for sparse sequential task, as it cannot recall visited places or task-relevant resources seen in the past. To address this, we propose MrSteve which stores novel events from visited places for efficient sparse sequential task-solving.
@@ 3.1 Memory Module: Construction of Place Event Memory
Specifically, each place cluster $M_{k}$ is subdivided into event clusters, denoted as $\smash{\{E_{i}^{k}\}_{i=1}^{d_{k}}}$, where each $\smash{E_{i}^{k}}$ represents the $i$-th event cluster in $k$-th place cluster, characterized by a center embedding $e_{E_{i}}^{k}$. These event clusters are newly created and updated as the place cluster accumulates a certain number of additional experience frames. For generating event clusters, DP-Means algorithm (Dinari & Freifeld, 2022) is applied on video embeddings of these frames, generated by MineCLIP, and the resulting cluster centers become a center embedding of each cluster. If the cosine similarity of the cluster centers between a newly created cluster and an existing cluster is higher than threshold $c$ (we define that two event clusters are indistinct), the two clusters are merged to prevent redundancy and ensure distinct event clusters within each place cluster. When memory capacity is exceeded, the oldest frame in the largest event cluster is removed, thus the memory can retain diverse places and distinct events within each place. More details on Place Event Memory can be found in Appendix E.
@@ 4.3 Memory-Constrained Task Solving with Memory
We demonstrated in Section 4.2 that memory is essential in solving sparse sequential tasks when there is no limitation in memory capacity. However, in real-world scenarios where memory capacity is limited, memorizing visited places and novel events becomes important. In this section, we show that Place Event Memory can benefit in this scenario. To verify this, we introduce three Memory Tasks, which are Find Water, Find Zombies’ Death Spot, and Find First-Visited House. In all tasks, the agent begins with an exploration phase, followed by a task phase. In the exploration phase, the agent follows a fixed exploratory behavior. In the subsequent task phase, the agent is given a MineCLIP embedding $e_{t}=\text{Enc}_{v}(\{i_{t}\}_{n=1}^{16})$ as a task embedding instead of text instruction, where $i_{t}$ is the pixel observation seen in the exploration phase. Then, this becomes an image goal navigation task where an agent should navigate to the location of the given image.
@@ 6 Conclusion
In this paper, we introduced MrSteve (Memory Recall Steve), a novel low-level controller designed to address the limitations of current LLM-augmented hierarchical approaches in general-purpose embodied AI environments like Minecraft. We argued that the primary cause of failures in many low-level controllers is the absence of an episodic memory system. To overcome this, we equipped MrSteve with Place Event Memory (PEM), a form of episodic memory that captures and organizes what, where, and when information from episodes. This allows for efficient recall and navigation in long-horizon tasks, directly addressing the limitations of existing low-level controllers like Steve-1, which rely heavily on short-term memory. Additionally, we proposed an Exploration Strategy and a Memory-Augmented Task Solving Framework, enabling agents to effectively switch between exploration and task-solving based on recalled events. Our results demonstrate significant improvements in both task-solving and exploration efficiency compared to existing methods. We believe that MrSteve opens new avenues for improving low-level controllers in hierarchical planning and are releasing our code to facilitate further research in this field.
@@ Acknowledgement
This work was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165), Brain Pool Plus Program (No. 2021H1D3A2A03103645), and Basic Research Lab (No. RS-2024-00414822) through the National Research Foundation (NRF) funded by the Ministry of Science and ICT (MSIT). The authors would like to thank the members of Machine Learning and Mind Lab (MLML) for helpful comments.
@@ Appendix A Limitations of Memory System in LLM-Augmented Agents
Memory systems in agents primarily aim to retrieve robust and accurate high-level plans for long-horizon tasks (Wang et al., 2023a; b; Zhu et al., 2023; Qin et al., 2024; Li et al., 2024). Existing works use an abstracted memory that stores the succeeded task with its plans, and often with history of observations for reliable retrieval, which is helpful when the similar plan in other tasks already exists in memory. However, these types of memory systems are not well-suited for low-level controllers for the following reasons:
  • Issues with Managing Memory Recent memory systems in Minecraft, when saving successful plans or skills, the history of observations for solving the task is all stored in FIFO manner (Wang et al., 2023b) or only task-relevant frames are stored in the plan (Li et al., 2024). However, computation complexity for retrieving the experience frames that are relevant to a new task is computationally expensive or even impossible (because the latter may not store the experience frames despite the agent observing it).
  • Lack of Mechanism for Retrieving Experience frames Current memory systems store the text instructions of succeeded tasks as keys and their plans to complete those tasks as values. The retrieval process begins by matching a task query to the task keys in memory and then filtering further using the current scene’s similarity to the stored frames before retrieving the final plan. These memory systems are targeted to retrieve the succeeded plan, but they lack a mechanism for utilizing the experience frames in the memory, which could be crucial for future tasks.
@@ C.1 Environments Setting
All tasks are implemented using MineDojo (Fan et al., 2022b). We utilize MineDojo’s success checker, where the success of each task is determined based on changes in the agent’s inventory. Hence, the agent succeeds in the task if the corresponding target item appears in the inventory. If the agent exceeds the time limit of each task or dies before completing assigned tasks, indicating the agent failed the task.
@@ C.2.2 Exploration & Navigation Task Details
For this task, we used Map 1 in Figure 8 with $100\times 100$ size, surrounded by high walls. The map includes complex terrains such as mountains and a water pond, which requires a robust low-level navigation policy for successful exploration. For each episode, the agent spawns in the center of the map and explores for 6,000 steps.
@@ C.2.4 Memory Tasks
Find Zombies’ Death Spot Task At the beginning, the agent sees burning zombies for the first 1K steps. Approximately 500 steps after the start of the episode, the zombies disappear, resulting in the agent observing two distinct scenes in the same location. A goal image for navigation is selected from the observation where the zombies are burning. After 1K steps, the agent starts to travel for the rest of the exploration phase. Once the exploration phase finishes, the agent returns to the place where the zombies were burning.
@@ Appendix E Place Event Memory, Place Memory, Event Memory Details
We provide the Algorithms on Write & Read operations of Place Event Memory and other memory variants, which are Place Memory and Event Memory. We also provide the specifications of each memory in the following section E.4.
@@ E.1 Place Event Memory Write & Read Operations
Event Cluster Details We use the DP-Means algorithm for clustering the events. DP-Means algorithm is a Bayesian non-parametric extension of the K-means algorithm based on small variance asymptotic approximation of the Dirichlet Process Mixture Model. It doesn’t require prior knowledge on the number of clusters $K$. To run this algorithm, we first set the initial number of clusters $K^{\prime}$ and cluster the data with K-Means++ initialization ($K^{\prime}$ can be $1$), then DP-Means algorithm automatically re-adjust the number of clusters based on the data points and cluster penalty parameter $\delta$. Thus, DP-Means algorithm behaves similarly to K-means with the exception that a new cluster is formed whenever a data point is farther than $\delta$ away from every existing cluster centroid.
@@
Place Cluster Details We provide a detailed explanation of how place clusters are formed. Each place cluster stores the agent’s experience frames based on its 2D ground position $(x,y)$ and head direction angle yaw. The size of a place cluster and its yaw range are defined by parameters $C$ (in Minecraft block) and $W\in(0^{\circ},180^{\circ})$, respectively, which distinguish different clusters. A place cluster is centered at a specific position $(x,y)$ with a center yaw $w$. An experience frame is assigned to a place cluster if the agent’s position falls within the range $(x-C/2,x+C/2)$ for $x$ and $(y-C/2,y+C/2)$ for $y$, and the yaw angle satisfies $(w-W/2,w+W/2)$. We use relative position and yaw for center of the place cluster, implying that the center of the initial place cluster has position $(0,0)$, and yaw 00.
@@ E.3 Event Memory Write & Read Operations
Discussions on MineCLIP threshold As in Event Memory’s Write Operation in Algorithm 7, clusters generated by DP-Means algorithm are either merged with an existing event cluster, or added as new event cluster. This is determined by MineCLIP threshold $c$, which is the criterion for separating event clusters. We note that using proper value for threshold is important for Event Memory to work reasonably across the tasks. If $c$ is too small, Event Memory will cluster experience frames with only a few clusters, which may not be visually distinctive. If $c$ is too large, in extreme case, Event Memory will make event cluster for each experience frame. When memory capacity is exceeded, it will randomly remove the oldest frame (since all event clusters have the same size), which behaves as a FIFO Memory.
Drawback in Event Memory Event Memory removes the frame in the largest event cluster, when memory capacity is exceeded. This indicates that the memory can retain visually distinct events. However, since it does not consider position information in clustering, similar visual frames in different places can be clustered into the same event cluster. This can be fatal since task-relevant frames in different places can be removed from the memory, which can be crucial in upcoming tasks. This drawback is shown in Find First-Visited House task in Section 4.3, and Long-Instruction task in Section 4.4, where MrSteve-EM showed suboptimal performances.
@@ Appendix G Exploration Experiments Details
Since Steve-1 (Lifshitz et al., 2024) has no memory module, it tends to visit the same place multiple times. We also observed that Steve-1 exhibits a behavior of moving straight ahead when task instruction "Go Explore" is given, often colliding with walls and then following along the wall instead of avoiding it, which harms the Map Coverage and Revisit Count.
@@ Appendix I MrSteve in Randomly Generated Maps
In this experiment, task A could involve collecting either water, beef, or wool, while task B involves gathering logs. Additionally, we introduce a sequential task named SEQ(4), which requires the agent to solve four consecutive tasks: log, water, wool, and beef. For the ABA-Random tasks, the agent was given 12K steps, and for the SEQ(4) task, 16K steps were allowed. We evaluated the following agents: MrSteve, MrSteve-EM, MrSteve-PM, MrSteve-FM, PMC-MrSteve-FM, and Steve-1.
As shown in Figure 13, MrSteve and its memory variants consistently showed higher success rate than Steve-1 across all tasks. This is because when Steve-1 is finished with task B, it tries to solve the final task A from scratch, making it difficult to complete all tasks in a limited time. However, MrSteve could retain the experience frames of task A in memory, making it efficient to solve all tasks in time. This result suggests that augmenting memory plays a crucial role in solving sequential tasks, even in randomly generated map.
@@
In this section, we study the robustness of our proposed agent MrSteve by exploring the effects of top-$k$ selection in Mode Selector and the size of a place cluster for Place Event Memory. For this, we used one of the tasks from ABA-Sparse task in Section 4.2, which is Wool-Log-Wool task. The success rate of MrSteve with different top-$k$’s and place cluster sizes are illustrated in Figure 14. For top-$k$ experiment, we tested five $k$ values, which are $1$, $2$, $4$, $10$, and $30$. From the results, we found that using small $k$ did not largely affect the performance. For cluster size experiment, we tested five cluster sizes, which are $1$, $2$, $4$, $6$, and $12$ in Minecraft blocks. From the results, we see that using bigger cluster size slightly lowered the performance since the center embedding of the place cluster may not be good summarization vector of the place when cluster size is large, making the agent difficult to recall task-relevant frames in the memory.
@@
Table 4: Querying Time and FLOPs for each memory type and $k$ values. We report the averages and standard errors over 1K query operations.
@@
In the previous section, we discussed the naive way of combining the idea of goal-conditioned behavior cloning and RL finetuning for training goal-conditioned navigator. In this section, we conduct a model architecture search on VPT model to find the optimal goal-conditioned navigation model. To do this, we focused on three key design choices: 1) the input location of goal embedding, 2) training parameters, and 3) different Tr-XL adaptors. For the location of goal embedding, we consider the three tactics: (1) add the goal embedding to the image embeddings and input to Tr-XL (TrXL-Cond.), (2) add the goal embedding to Tr-XL output (Head-Cond.), and (3) add the goal embedding both at the input and output of Tr-XL (Dual-Cond.). For training parameters, we use full-finetuning or finetuning only part of the model. For the Tr-XL adaptors, we consider the adaptor from Houlsby et al. (2019), and LoRA adaptor (Hu et al., 2021b). In Figure 15, we illustrate the different model architectures for goal-conditioned VPT finetuning models.
For the model search, we summarized the navigation performance (SPL and Success Rate (SR)) of different model architectures in Table 6. Interestingly, we found that adding goal embedding to image embeddings (TrXL-Cond.) showed suboptimal behavior, while adding goal embedding to output of Tr-XL (Head-Cond., Dual-Cond.) showed good performance indicating that propagating learning signal to goal embedding through Tr-XL is difficult from RL objective. We note that using LoRA adaptor showed higher performance than adaptors from Houlsby et al. (2019). In conclusion, using Head-Cond. with LoRA finetuning performed the best. Additionally, we tested one additional model that does not update Tr-XL, which showed comparable result to the best performing model. This architecture benefits from using a smaller number of learning parameters and lower gradient computations. Thus, we use the VPT-Nav trained with one of two architectures, which are Head-Cond. with LoRA finetuning and Head-Cond. with No Tr-XL Update.
@@
Table 8: Long-Horizon Planning Tasks details. The required subgoals denote the length of the shortest plan for each task. The items that the low-level controller should obtain are listed in Items to Obtain column.
@@
LLM-generated initial plans can be inaccurate, so the low-level controllers often fail to execute subgoals. For instance, we found that the initial plans frequently omit the subgoal for creating crafting table, which is prerequisite item for the tasks in Table 8 except the carpet task. To address this problem, DEPS framework introduced the replanning procedure with the descriptor and explainer modules. The descriptor module makes a text prompt representing the inventory of the agent. The explainer modules ask the LLM the reason of failure based on the text prompt from the descriptor module. Based on the explanation, the high-level planner asks the LLM to revise the plan and the low-level controller executes subgoals from the revised plan.
Additionally, because the LLM does not observe the environment, the order of subgoals in a plan can be suboptimal. When some subgoals share the same prerequisite so they can be executed in any order, it is more efficient to execute subgoal, which can be completed faster than the other subgoals, first. This can prevent wasting time by giving up subgoals that could be completed quickly, pursuing a subgoal whose completion time is uncertain, and then going back to search for the previously quicker task again after executing the subgoal. In DEPS, this procedure called elector module is implemented with the horizon prediction module in GSB (Cai et al., 2023a). The horizon prediction module $\mu(s_{t},\tau_{t,i}^{g})$ takes the current observation $s_{t}$ and the textual subgoals $\{\tau_{t,i}^{g}\}_{i=1}^{N}$ and predicts the time to complete each subgoal. Based on the time prediction, a subgoal to be executed is sampled from the distribution as follows: